### v1

Abstract
While large language models (LLMs) achieve near-perfect scores on medical licensing exams, these evaluations inadequately reflect the complexity and diversity of real-world clinical practice. We introduce MedHELM, an extensible evaluation framework for assessing LLM performance for medical tasks with three key contributions. First, a clinician-validated taxonomy spanning 5 categories, 22 subcategories, and 121 tasks developed with 29 clinicians. Second, a comprehensive benchmark suite comprising 35 benchmarks (17 existing, 18 newly formulated) providing complete coverage of all categories and subcategories in the taxonomy. Third, a systematic comparison of LLMs with improved evaluation methods (using an LLM-jury) and a cost-performance analysis. Evaluation of 9 frontier LLMs, using the 35 benchmarks, revealed significant performance variation. Advanced reasoning models (DeepSeek R1: 66% win-rate; o3-mini: 64% win-rate) demonstrated superior performance, though Claude 3.5 Sonnet achieved comparable results at 40% lower estimated computational cost. On a normalized accuracy scale (0-1), most models performed strongly in Clinical Note Generation (0.73-0.85) and Patient Communication & Education (0.78-0.83), moderately in Medical Research Assistance (0.65-0.75), and generally lower in Clinical Decision Support (0.56-0.72) and Administration & Workflow (0.53-0.63). Our LLM-jury evaluation method achieved good agreement with clinician ratings (ICC = 0.47), surpassing both average clinician–clinician agreement (ICC = 0.43) and automated baselines including ROUGE-L (0.36) and BERTScore-F1 (0.44). Claude 3.5 Sonnet achieved comparable performance to top models at lower estimated cost. These findings highlight the importance of real-world, task-specific evaluation for medical use of LLMs and provides an open source framework to enable this.
keywords:
large language models, evaluation, medicine, benchmark, taxonomy
1 Introduction
Large Language Models (LLMs) have shown impressive performance on medical knowledge benchmarks, achieving $\sim$99accuracy on standardized exams like MedQA [1]. This has sparked interest in deploying them in healthcare settings: supporting clinical decision-making such as diagnosis and treatment [2], optimizing clinical workflows including documentation and scheduling [3], and enhancing patient education and communication [4].
However, there is a large gap between performance on medical knowledge benchmarks and readiness for real-world deployment due to three key limitations in these existing benchmarks: (1) Questions do not match real-world settings – Existing benchmarks rely on synthetic vignettes or narrowly-scoped exam questions, failing to capture key aspects of real diagnostic processes such as extracting relevant details from patient records [5, 6]. (2) Limited use of real-world data – Only 5% of LLM evaluations use real-world electronic health record (EHR) data [7]. EHRs contain ambiguities, inconsistencies, and domain-specific shorthand that synthetic data cannot replicate. (3) Limited task diversity – Around 64% of LLM evaluations in healthcare focus only on medical licensing exams and diagnostic tasks [7], ignoring essential hospital operations such as administrative tasks (e.g., generating prior authorization letters, identifying billing codes), clinical documentation (e.g., writing progress notes or discharge instructions), and patient communication (e.g., asynchronous messaging through electronic patient portals) [8].
Recent work on HealthBench [9] has advanced the evaluation of LLMs in medicine by scoring 5000 single-turn, free-text dialogues in which the model acts independently, much like a direct-to-patient advice line, without follow-up questions or human oversight. Its physician-authored rubrics reward exhaustive, risk-averse responses that maximize safety and completeness, offering a valuable stress test for fully autonomous chatbots. However, this design does not capture the iterative, context-aware interactions clinicians expect from an assistive co-pilot, nor does it assess performance on structured tasks that dominate everyday workflows (e.g., order review, note generation, literature summarization).
To address these limitations, we introduce MedHELM (Holistic Evaluation of Large Language Models for Medical Tasks), an extensible evaluation framework for assessing LLM performance in completing medical tasks (Figure 1). Inspired by the HELM project’s standardized cross-domain evaluations [10], using MedHELM  we evaluate 9 LLMs using 35 distinct benchmarks covering all 22 subcategories of medical tasks, focusing on clinicians’ day-to-day activities beyond just taking licensing exams. We assess performance using benchmark-appropriate metrics (i.e., exact match for closed-ended benchmarks, LLM-jury where three LLMs evaluate responses using tailored rubrics for open-ended benchmarks with demonstrated agreement to clinician ratings) as well as estimated computational cost to provide practical deployment insights. Our primary contributions are:
  1. 1.
    Clinician-validated taxonomy: A 5-category, 22-subcategory, 121-task taxonomy developed with 29 clinicians. Clinicians achieved a 96.7% agreement rate when mapping subcategories to appropriate categories, validating the clear and discrete nature of the taxonomy. We present the complete taxonomy in the Results and Appendix sections.
  2. 2.
    A benchmark suite with full taxonomy coverage: A collection of 35 benchmarks spanning all 22 subcategories of medical tasks. This includes 17 existing benchmarks, 5 re-formulated benchmarks based on existing datasets, and 13 new benchmarks. [^1] For privacy and regulatory compliance, as well as to prevent inclusion in LLM training data, 14 datasets are not publicly released.
  3. 3.
    Comparative evaluation of models along with cost-performance analysis: A systematic evaluation shows that reasoning models achieve the highest overall performance. For example, DeepSeek R1 and o3-mini score the best in Clinical Note Generation (0.85) while o3-mini leads in Clinical Decision Support (0.72) and Medical Research Assistance (0.75)  [^2] Note: Values in parentheses represent scores on a normalized accuracy scale (0-1).. On a subset of samples from two open-ended benchmarks (MEDIQA-QA & ACI-Bench), our LLM-jury approach achieves 0.47 agreement (as measured by intraclass correlation coefficient) with clinicians, beating average clinician-to-clinician agreement (0.43) and outperforming ROUGE-L (text overlap metric, 0.36) and BERTScore-F1 (semantic similarity metric, 0.44). A cost-performance analysis identifies Claude 3.5 Sonnet as providing comparable results with lower estimated computational cost.
The MedHELM framework addresses a critical need in the testing and evaluation of AI for medical use by providing consistent, real-world evaluation standards for medical (and healthcare) applications of LLMs. This framework benefits three key stakeholder groups: (1) Healthcare systems evaluating LLMs for certain tasks, (2) AI developers identifying performance gaps across medical tasks, and (3) Researchers developing methods to reproducibly measure LLM capabilities on medical tasks. To foster collaborative improvement of such AI evaluation, we provide an openly accessible leaderboard with current benchmarking results and share the codebase  [^3] https://github.com/stanford-crfm/helm, with documentation  [^4] https://crfm-helm.readthedocs.io/en/latest/medhelm/ for contributing new datasets, evaluation metrics, and benchmarking custom models. By standardizing terminology and evaluation methods across the task taxonomy, MedHELM establishes a foundation for reproducible and real-world assessment of AI capabilities in medicine.
Figure 1: This figure illustrates: (a) a clinician-validated taxonomy organizing 121 medical tasks into 5 categories and 22 subcategories; (b) a suite of benchmarks that map existing benchmarks to this taxonomy and introduces new benchmarks for complete coverage; and (c) an evaluation comparing reasoning and non-reasoning LLMs, with model rankings, LLM jury based evaluation of open-ended benchmarks, and cost-performance analysis
2 Results
2.1 Clinician Validation of the Taxonomy
In a structured review process, clinicians evaluated the initial taxonomy comprising 5 categories, 21 subcategories, and 98 tasks. When asked to assign each subcategory to its appropriate top-level category, clinicians correctly matched 96.7% of subcategories to their intended categories. The clinicians rated the comprehensiveness of the proposed tasks at a mean of 4.21/5 (n = 29) and provided 107 comments with suggestions for improvement. Based on this feedback, we refined task definitions and expanded the taxonomy to its final form: 5 categories, 22 subcategories, and 121 tasks. An overview of the final taxonomy is shown in Figure 2, with the complete list provided in Appendix A.
Figure 2: Overview of the final taxonomy comprising 5 main categories and 22 subcategories.
2.2 Overview of the Benchmark Suite
Our 35 benchmarks span all 22 subcategories, providing full coverage over categories and subcategories in our taxonomy (Table 7). The benchmark suite comprises 17 existing benchmarks, 5 re-formulated benchmarks derived from previously unevaluated medical datasets, and 13 new benchmarks, out of which 12 are EHR-based. The suite includes 13 open-ended benchmarks (requiring free-text generation) and 22 closed-ended benchmarks (with predefined answer choices). Access levels are designated as 14 public, 7 gated (i.e. requiring approval) and 14 private.
Clinical Decision Support is the most represented category with 10 benchmarks, followed by Patient Communication (8), Clinical Note Generation and Medical Research Assistance (6), and Administration & Workflow (5). The distribution of benchmarks across subcategories is uneven, with 15 subcategories containing a single benchmark and the remaining 7 subcategories containing between 2 and 5 benchmarks each.
2.3 Model Evaluation & Cost-Performance Analysis
2.3.1 Overall Performance
Pairwise Win-Rate and Average Scores
Table 1 compares models using win-rate and macro-average performance metrics (defined in the table caption). DeepSeek R1 performed best, winning 66% of head-to-head comparisons with a macro-average of 0.76 and low win standard deviation (0.10). o3-mini followed with a 64% win rate and the highest macro-average (0.78), driven by strong performance in benchmarks in the clinical decision support category.
The Claude models achieved 63-64% win rates and identical macro-averages (0.74). GPT-4o achieved a 57% win rate, while Gemini 2.0 Flash (42%) and GPT-4o mini (39%) performed lower. Open-source Llama 3.3 Instruct achieved a 30% win rate. Gemini 1.5 Pro ranked lowest with 24% wins but had the lowest win standard deviation (0.08), showing the most consistent competitive performance.
[…]
DeepSeek R1 0.66 0.10 0.76 0.22
o3-mini (2025-01-31) 0.64 0.16 0.78 0.19
Claude 3.7 Sonnet (20250219) 0.64 0.13 0.74 0.21
Claude 3.5 Sonnet (20241022) 0.63 0.14 0.74 0.21
GPT-4o (2024-05-13) 0.57 0.17 0.75 0.19
Gemini 2.0 Flash 0.42 0.17 0.72 0.21
GPT-4o mini (2024-07-18) 0.39 0.18 0.73 0.20
Llama 3.3 Instruct (70B) 0.30 0.13 0.71 0.22
Gemini 1.5 Pro (001) 0.24 0.08 0.69 0.21
Table 1: Comparison of performance of frontier models across 35 MedHELM benchmarks, sorted by descending win-rate. Bold indicates the best value in each column. Win-rate represents the proportion of pairwise comparisons where each model achieved superior performance across all 35 benchmarks (possible range: 0-1). Win standard deviation (SD) measures how consistently a model wins (lower values = more consistent). Macro-avg is the average performance score across all 35 benchmarks. SD shows how much performance varies across different benchmarks (lower values = more consistent across benchmarks).
Performance by Benchmark
We present every model’s normalized score from each of the 35 benchmarks as a heatmap in Figure 3, where darker green indicates higher performance. We see that models perform worse on benchmarks such as MedCalc-Bench (calculating medical values from patient notes), EHRSQL (generating SQL queries from natural language instructions for clinical research—originally intended as code generation dataset), and MIMIC-IV Billing Code (assigning ICD-10 codes to clinical notes). The best performance is seen for the NoteExtract benchmark (extracting specific information from clinical notes). Broader category-level trends are described below.
Figure 3: Heatmap of normalized scores (0–1) for each model (rows) across 35 benchmarks (columns). Dark green indicates high performance; dark red indicates low performance. The statistical significance of model differences varies by benchmark, with an analysis of minimum detectable effects shown in Appendix E. Metrics include EM as exact match, Jury Score as the average normalized score from 3 frontier LLMs, MedFlagAcc as binary accuracy for detecting presence of errors, MedCalcAcc as exact match or thresholded match depending on question type, and EHRSQLExecAcc as execution accuracy of generated code against a target output.
Performance by Category
Figure 4 presents performance scores by the 5 top-level categories in the taxonomy. Most models achieve their highest mean scores in Clinical Note Generation (0.73-0.85, with Gemini 1.5 Pro as an outlier at 0.56) and Patient Communication & Education (0.78-0.83), moderate performance in Medical Research Assistance (0.65-0.75), and generally lower scores in Clinical Decision Support (0.56-0.72) and Administration & Workflow (0.53-0.63). Performance variations across categories reflect distinct task-based challenges: free-text generation tasks (Clinical Note Generation, Patient Communication) leverage models’ natural language strengths, while structured reasoning tasks (Clinical Decision Support, Administration) require domain-specific knowledge integration and logical inference. These patterns have important implications for selective deployment strategies in healthcare settings.
At the model level, DeepSeek R1 and o3-mini lead across most categories, with DeepSeek R1 excelling in Clinical Note Generation (0.85, tied with o3-mini) and Patient Communication & Education (0.82), while o3-mini leads in Clinical Decision Support (0.72) and Medical Research Assistance (0.75). GPT-4o achieves the highest score in Patient Communication & Education (0.83). The Claude Sonnet series (3.5 and 3.7) shows consistent performance across categories, with scores of 0.78-0.81 in Clinical Note Generation and Patient Communication & Education. The Gemini series of models shows significant performance variability across categories, performing well in Patient Communication & Education tasks (0.80) but poorly in other categories. This likely reflects stricter safety guardrails leading to more question refusals. The open-source Llama 3.3 performs well in Patient Communication & Education (0.82) but shows the lowest score in Administration & Workflow (0.53), indicating areas for future improvement.
Figure 4: Mean normalized scores (0-1 scale) across the 5 categories for all evaluated models. Darker green represents higher scores. Models are ordered by mean win rate from top (highest) to bottom (lowest), while categories are arranged left to right.
2.3.2 Evaluation of Open-Ended Benchmarks
For our 13 open-ended benchmarks, we implemented an LLM-jury evaluation approach. To assess this method’s validity, we collected independent clinician ratings on a subset of model outputs. We used 31 instances from ACI-Bench and 25 from MEDIQA-QA to compare clinician-assigned scores to the jury’s aggregated ratings as described in the methods. Table 2 reports intraclass correlation coefficient (ICC(3,k)) values (after rater-wise $z$-scoring) for the LLM-jury versus clinicians, alongside ROUGE-L (text overlap metric) and BERTScore-F (semantic similarity metric) baselines, and the average clinician–clinician agreement.
Overall, the LLM-jury achieves an ICC of 0.47, which beats the average clinician–clinician agreement (ICC = 0.43) and automated baselines including ROUGE-L (0.36) and BERTScore-F (0.44). These results demonstrate that our LLM-jury mirrors clinician judgment better than standard lexical metrics, establishing its validity as a stand-in for clinician raters.
Benchmark LLM ROUGE-L BERTScore-F Clinician
Combined 0.474 (0.100, 0.690) 0.361 (0, 0.630) 0.441 (0.050, 0.670) 0.426 (0.295, 0.585)
ACI-Bench 0.305 (0, 0.670) 0.445 (0, 0.730) 0.250 (0, 0.640) 0.458 (0.201, 0.945)
MEDIQA 0.625 (0.150, 0.830) 0.343 (0, 0.710) 0.668 (0.250, 0.850) 0.520 (0.500, 0.534)
Table 2: Agreement of LLM-jury and of automated metrics with clinician ratings The table entries are ICC(3,$k$) coefficients after $z$-scoring within rater (ICC3k-$z$); 95 % confidence intervals are shown in parentheses. Higher ICC indicates better agreement with clinician ratings. The last column gives the average clinician–clinician agreement for each dataset.
2.3.3 Cost–Performance Analysis
We estimated the cost (USD) of evaluating each model based on publicly listed pricing as of 05/12/2025, using the total input tokens and maximum output tokens consumed during benchmark runs and LLM-jury evaluation. These costs are an upper-bound estimate, since models may generate fewer tokens than the maximum allowed output. We plotted the mean win-rate against the cost (Figure 5, with a detailed breakdown in Table 3, and inference costs in Appendix G). As expected, non-reasoning models—GPT-4o mini (\$805) and Gemini 2.0 Flash (\$815)—incurred the lowest costs and achieved win-rates of 0.39 and 0.42, respectively. Open-source Llama 3.3 Instruct (\$940) had a 0.30 win-rate, while Gemini 1.5 Pro (\$1,130) reached 0.24. Reasoning models incurred higher costs, DeepSeek R1 (\$1,806) and o3-mini (\$1,722), with win-rates of 0.66 and 0.64 respectively. Claude 3.5 Sonnet (\$1,571) and Claude 3.7 Sonnet (\$1,537) provide a good cost-performance balance, achieving $\sim 0.63$ win-rate at reduced costs.
Figure 5: Scatter plot of mean win-rate (y-axis) versus estimated computational cost (x-axis) for each of the 9 models across 35 benchmarks. Each point represents a model, with the position indicating the relationship between performance (y-axis) and total cost of evaluation including benchmark runs and evaluation by LLM-jury (x-axis). Costs represent upper-bound estimates based on maximum output token usage.
Model Model Creator Window Size Access Benchmark Tokens Jury Tokens Benchmark Cost Jury Cost Total Cost
Claude 3.5 Sonnet (20241022) Anthropic 200,000 Closed 245,958,343 45,868,483 \$778.67 \$792.21 \$1,570.88
Claude 3.7 Sonnet (20250219) Anthropic 200,000 Closed 242,510,517 42,326,449 \$768.26 \$768.38 \$1,536.64
Gemini 1.5 Pro (001) Google 1,000,000 Closed 277,273,571 42,864,352 \$359.33 \$771.73 \$1,131.06
Gemini 2.0 Flash Google 1,000,000 Closed 276,777,154 42,864,352 \$43.04 \$771.73 \$814.77
GPT-4o (2024-05-13) OpenAI 128,000 Closed 248,731,118 41,929,516 \$647.28 \$765.91 \$1,413.19
GPT-4o mini (2024-07-18) OpenAI 128,000 Closed 248,731,118 41,929,516 \$38.83 \$765.91 \$804.74
Llama 3.3 Instruct (70B) Meta AI 128,000 Open 242,895,608 42,121,933 \$172.45 \$767.13 \$939.58
DeepSeek R1 DeepSeek 128,000 Open 334,133,126 68,364,208 \$848.21 \$957.96 \$1,806.17
o3-mini (2025-01-31) OpenAI 128,000 Closed 337,967,189 68,617,978 \$762.50 \$959.54 \$1,722.04
Table 3: Comparison of LLMs by architecture, access type, and token usage metrics across MedHELM. Benchmark tokens denote total input/output tokens in completing the medical task represented by the benchmark; jury tokens are tokens used for open-ended evaluations via the LLM-jury. Total cost reflects the estimated per-model expense of running a MedHELM evaluation across 35 benchmarks and represents an upper bound based on maximum output token usage. For 9 models and 35 benchmarks in the current MedHELM suite, one update to the leaderboard is estimated to cost $11,739.07$.
3 Discussion
We present a framework for assessing LLM performance for real-world medical tasks. Our clinician-validated taxonomy provides a structure for summarizing models’ strengths and limitations across medical tasks. High clinician agreement (96.7%) in assigning subcategories to categories suggests the taxonomy effectively captures how healthcare professionals conceptualize their work. Our benchmark suite reveals nuances in model capabilities that are not seen with medical knowledge benchmarks alone. The higher performance in communication tasks than administrative ones may stem from administrative workflows using data not seen during training, warranting caution in healthcare AI implementation for back-office tasks without quantifying task-specific performance. This framework addresses the primary limitations of current benchmarks, such as the lack of using real-world data, use of evaluation setups that do not match real-world settings, and limited task diversity. The cost-performance trade-off shows that while reasoning models have superior performance, their substantially higher costs may not justify their deployment for all tasks. In a resource-constrained setting, models such as Claude 3.5 Sonnet offer a balance, achieving a win-rate of 0.63 at a lower cost. Finally, our LLM-jury based approach addresses a critical gap in current evaluation approaches. By beating clinician-clinician agreement, this approach enables scalable evaluation of open-ended model outputs without requiring extensive clinician time — a scarce and expensive resource.
Several limitations remain. While our LLM-jury approach was validated on only two benchmarks, expanding clinician annotations across more benchmarks would strengthen the clinician–LLM agreement estimates. Our current rubrics operate at the benchmark level, but instance-level rubrics could provide better evaluation, particularly for subjective or context-dependent medical tasks where gold standard responses may not exist. Such approaches would further scale LLM-jury evaluation beyond reliance on gold standard responses [11]. Administration & Workflow emerged as the weakest performance area for all models. Understanding the underlying causes of this poor performance, whether stemming from training data limitations, task complexity, or distributional shifts, is essential for safe deployment in healthcare operations.
In conclusion, MedHELM provides a comprehensive framework for assessing LLM performance across real-world medical tasks through our clinician-validated taxonomy (96.7% agreement) spanning 5 categories, 22 subcategories, and 121 tasks. Our benchmark suite of 35 datasets reveals that most models perform best in Clinical Note Generation (0.73-0.85, with one outlier at 0.56) and Patient Communication & Education (0.78-0.83), moderately in Medical Research Assistance (0.65-0.75), and worst in Clinical Decision Support (0.56-0.72) and Administration & Workflow (0.53-0.63). Reasoning models DeepSeek R1 and o3-mini led overall with win-rates of 0.66 and 0.64 respectively, though Claude models offer competitive performance (win-rate of $\sim 0.63$) at lower computational cost. Through our public leaderboard [^5] https://crfm.stanford.edu/helm/medhelm/latest/ and shared codebase [^6] https://github.com/stanford-crfm/helm, MedHELM establishes infrastructure for ongoing collaborative assessment as models evolve, advancing medical AI evaluation that better reflects the complexity of real-world medical practice.
4 Methods
Motivation and related work. Most LLM evaluations in medicine still rely on closed-form question-answering over exam-style datasets such as MedQA and MedMCQA, with only $\sim$5incorporating real EHR data and very few addressing free-text generation tasks or cost-aware metrics [12, 6, 13]. Large-scale Natural Language Processing (NLP) meta-benchmarks (HELM, BIG-Bench) demonstrate the value of task diversity and multi-metric scoring [10, 14], while biomedical efforts like ClinicBench [15] or MMedBench [16]) each advance a single dimension (e.g. multimodality or cost-aware metrics) without clinician-validated scope or extensible tooling. Recent work on HealthBench [9] addresses some of these limitations by incorporating physician-developed rubrics for 5,000 health conversations, but lacks a comprehensive taxonomy of medical tasks and is done with synthetic data. Our goal is to close these gaps by co-designing, with clinicians, a taxonomy-guided benchmark suite that (i) spans the full spectrum of real medical work, (ii) uses both public and private medical record data, and (iii) uses evaluation protocols that align with clinician judgment.
4.1 Development of the Taxonomy of Tasks
Early attempts (BigBIO, ClinicBench) at creating a taxonomy of tasks in medical settings either harmonize datasets into broad categories without clinician input or collapse heterogeneous skills under a single “generation” label [17, 15]. To ensure our benchmarks accurately reflect the complexity of medical practice, we developed a taxonomy that mirrors how clinicians conceptualize their daily work. By defining a hierarchical structure, we guarantee that each benchmark maps to a concrete medical activity. Our taxonomy consists of three levels:
  • Category: A broad domain of medical activity (e.g., Clinical Decision Support).
  • Subcategory: A group of related tasks within a Category (e.g., Supporting Diagnostic Decisions).
  • Task: A discrete action taken during the delivery of medical care (e.g., Generate differential diagnoses).
This structure enables systematic coverage of the medical care landscape while maintaining clear boundaries between distinct activities that care providers perform.
4.1.1 Initial drafting
We based our taxonomy on tasks identified in a JAMA review [7]. Working with a clinician (MK), we reorganized these tasks into functional themes that reflect real-world activities, resulting in 98 distinct tasks organized into 21 subcategories within 5 categories:
  1. 1.
    Clinical Decision Support
  2. 2.
    Clinical Documentation
  3. 3.
    Patient Communication & Education
  4. 4.
    Medical Research Assistance
  5. 5.
    Administration & Workflow
Two principles guided our taxonomy development:
  • Medical relevance: Each task maps directly to actions routinely performed by care providers.
  • Clear boundaries: Categories and subcategories were defined to minimize overlap while preserving meaningful functional distinctions.
4.1.2 Validation
To validate our initial taxonomy, we designed a two-part survey completed by 29 practicing clinicians across 14 medical specialties. More information on participating clinicians can be found in Appendix B. The survey assessed both categorical organization and real-world relevance:
In the first section, clinicians assigned each of our 21 subcategories to one of the 5 main categories. This exercise tested whether our taxonomy structure matched how clinicians naturally organize medical tasks.
In the second section, clinicians evaluated the comprehensiveness of our taxonomy on a 5-point scale, where a score of 5 means that our categorization covered all routine medical tasks and a score of 1 means it covered very few. They also provided feedback through an open dialogue box where they could suggest missing tasks and recommend terminology improvements. This systematic validation approach evaluated both the taxonomy’s organizational logic and its comprehensiveness in representing actual medical tasks.
Based on the comments, we refined definitions and expanded the taxonomy to have 5 categories, 22 subcategories, and 121 tasks.
4.2 Construction of the Benchmark Suite
4.2.1 Curation of datasets
To construct a comprehensive suite of 35 benchmarks spanning our taxonomy, we employed a three-tiered dataset curation strategy:
  1. 1.
    Existing benchmarks: We incorporated existing benchmarks from public or gated sources (e.g., MedQA, MIMIC-IV Billing Code, ACI-Bench) to ensure broad subcategory coverage.
  2. 2.
    Reformulated benchmarks: We transformed previously unevaluated medical data collections into "reformulated benchmarks" by applying standardized prompt templates and specifying evaluation metrics. This approach addressed subcategories where datasets existed but lacked LLM-ready evaluation benchmarks.
  3. 3.
    New benchmarks: To address the under-representation of Administration & Workflow, we partnered with Stanford Healthcare to curate private datasets for tasks that are routinely done in health systems but for which benchmark datasets do not exist (e.g.  referral triage, scheduling).
Each benchmark is labeled by source type (existing / reformulated / new) and access level (public / gated / private), with the provenance documented in the MedHELM repository.
4.2.2 Specification of Prompts and Metrics
To transform each curated dataset into a MedHELM benchmark, we defined three components for every item in the dataset:
  • Context: the raw input presented to the LLM (e.g.  a clinical note, patient message, or structured EHR record).
  • Prompt: a standardized instruction template to elicit consistent, task-appropriate responses (e.g.  “Answer in 2–3 sentences” for open-ended summaries, or MCQ framing for multiple-choice questions).
  • Evaluation Metric: a pre-specified scoring method matched to the task type:
    • Exact-match accuracy for single-token or numeric outputs (e.g.  selecting the correct option in MedQA).
    • Micro-F1 for multi-label classification tasks (e.g.  ICD-10 code assignment).
    • LLM-jury ensemble for open-ended text generation: we use a three-model Likert-scale protocol assessing medical accuracy, completeness and clarity, and secondary metrics, ROUGE and BERTScore, to capture lexical and semantic overlap.
  • Gold Standard Response (Optional): the reference output (numeric result, classification label, or sample text) against which the model’s response is scored (e.g., "4" in response to "What’s a patient’s HAS-BLED score?"). While most benchmarks include a gold standard, our framework accommodates benchmarks without one, such as NoteExtract, providing flexibility for future evaluation needs.
Table 4 illustrates these components for the HeadQA benchmark (subcategory: Providing Clinical Knowledge Support) and the MedDialog benchmark (subcategory: Patient-Provider Messaging).
We found quality issues in some of the gold standard responses for reformulated benchmarks. For example, gold standard responses in the MIMIC-RRS benchmark occasionally contained information from a patient’s EHR that was not passed into the model’s context. To assess the impact of these low-quality gold standard responses, we conducted a sensitivity analysis by filtering "problematic" gold standard responses using an LLM judge and recalculating metrics (Appendix D). Model rankings remained unchanged, as instances with "problematic" and "non-problematic" gold standard responses received similar jury scores. This stability exists because our LLM-jury (prompt in Appendix F) uses gold standard responses only when needed.
Benchmark Context Prompt Evaluation Metric
HeadQA Question: In the stools of a 6-year-old boy with diarrhea and abdominal pain, a packet of parasite eggs surrounded by a transparent membrane is observed. Which treatment do you consider most appropriate? A. Praziquantel B. Azithromycin C. Metronidazole D. Albendazole You are a highly knowledgeable AI assistant specializing in biomedical sciences. Your task is to answer multiple-choice questions accurately based on the options provided. Select the correct answer by outputting only the letter corresponding to your choice (A, B, C, or D). Question: {Context} Answer: Exact Match Answer: D
MedDialog Patient: I have had this pain in my left arm down to almost my wrist. It first started where I would have a sharp pain between my armpit and boob. It doesn’t hurt all the time but has become more often lately. It’s an aching throbbing feeling but also a numbing weak feeling to where I lose my strength in my arm. After a while it will pass but I don’t know if it’s something serious. Need help. Doctor: Hi, take this seriously, it could be a lump compressing on your nerves, or any infection. You need an examination by a doctor to exactly judge what your symptoms mean and what is their cause. I recommend you consult your doctor at the earliest. Generate a one sentence summary of this patient-doctor conversation. LLM Jury Accuracy: 4.88
Table 4: Case study of HeadQA and MedDialog, two benchmarks under MedHELM
4.3 Model Evaluation and Cost-Performance Analysis
4.3.1 Model Selection and Inference Pipeline
We evaluated 9 state-of-the-art LLMs (Table 1) under a uniform prompting and decoding regimen. All models were queried via their respective APIs or local endpoints with sampling temperature set to 0 for deterministic outputs. All experiments were conducted on a PHI-compliant shared cluster maintaining full HIPAA compliance.
4.3.2 Performance Metrics
To quantify task performance, we computed:
  • Pairwise win-rate: For each of the 35 benchmarks, we compared each model against every other; a “win” is assigned if a model’s normalized score $\geq$ its rival’s. We then averaged wins over all pairings.
  • Macro-average score: The overall performance score calculated by averaging results across all 35 benchmarks, with each benchmark weighted equally regardless of size (0–1 scale).
4.3.3 Evaluation of Open-Ended Benchmarks
For open-ended benchmarks, early pipelines used a single “LLM-as-Judge” to score outputs [18, 19], but high variance and bias spurred the “LLM-as-Jury” paradigm, aggregating $k$ independent judgments for closer expert agreement [20, 21]. Extensions like G-Eval introduce chain-of-thought (CoT) self-critique per juror, and tools such as SelfCheckGPT and FActScore target hallucination and factuality [22, 23, 24]. MedHELM adopts a three-member jury without CoT to balance reliability and runtime.
Evaluation using LLM-Jury
We selected a three-member jury based on prior research demonstrating that odd-numbered panels reduce tie scenarios while maintaining reliability [25]. The jury composition (GPT-4o, Claude 3.7 Sonnet, LLaMA 3.3 70B) was chosen to represent diverse model architectures and training approaches, minimizing systematic bias from any single provider. We prompted each judge to score the model-generated responses on a 1–5 Likert scale according to three axes adopted from [26]:
  • Accuracy: Factual correctness and adherence to medical guidelines.
  • Completeness: Thoroughness in addressing all aspects of the query.
  • Clarity: Organization, readability, and easy to understand language.
For the NoteExtract benchmark, which requires restructuring free-form clinical care plans into a specified format without a gold standard response, we modified our evaluation approach. We replaced the "completeness" criterion with "structure," which assessed whether model responses properly reorganized the input text according to the requested format. The final LLM-jury score for each response is the mean of all nine ratings (3 judges × 3 axes).
Clinician Rating
To validate the LLM-jury approach, we collected clinician ratings on a subset of two open-ended benchmarks (MEDIQA and ACI-Bench). We selected these benchmarks because they were publicly available (facilitating annotation) and represent different categories (Patient Communication and Education and Clinical Note Generation, respectively). 20 clinicians across various specialties each scored a subset of responses on the same three axes, with at least two clinicians per instance.
Clinician–LLM Agreement Metrics
We assessed agreement between the LLM-jury and clinicians via two intraclass correlation coefficients (ICC), after applying z-score normalization to each rater’s composite (mean of the three axis ratings) to remove scale biases:
$z_{ij}=\frac{x_{ij}-\bar{x}_{i}}{s_{i}}\,,$
where $x_{ij}$ is rater $i$’s composite on instance $j$.
  1. 1.
    ICC(3,k)${}_{\text{z}}$: We treat the clinicians’ mean and LLM jury mean as two fixed raters in a two-way mixed-effects model, thus, we compute
    $\mathrm{ICC}(3,k)=\frac{MS_{\text{cases}}-MS_{\varepsilon}}{MS_{\text{cases}}},\quad k=2,$
    to quantify absolute agreement.
  2. 2.
[…]
    Together, these metrics measure (i) the fidelity of LLM-jury to expert clinician rating and (ii) whether LLM–Clinician alignment approaches inter-clinician agreement.
ROUGE and BERTScore were also computed for all open-ended benchmarks with gold standard responses but used only as secondary metrics due to their limited alignment with clinical judgment.
4.3.4 Cost–Performance Analysis
We tracked token usage for both benchmark prompts and jury evaluations across all models by counting input tokens and assuming the maximum allowed output tokens. This approach was necessary because, for o3-mini, it is not possible to fully observe the token usage as the thinking tokens don’t appear in the final output. As a result, our cost estimates represent an upper bound. Using published per-1M-token pricing as of 05/12/2025, we estimated total cost per model. Cost–performance trade-offs were visualized by plotting mean win-rate against estimated total evaluation cost, highlighting models that optimize accuracy per dollar spent.
By integrating clinician-validated taxonomy design, a balanced mix of public and private datasets, and LLM-jury evaluation, our benchmark addresses the three structural gaps identified in earlier work—task diversity, medical grounding, and metric fidelity—while providing transparent cost accounting for real-world deployment decisions.
Appendix A Task Taxonomy
  1. 1.
    Overview 1
  2. 2.
    Clinical Decision Support 2
    • Definition
    • Subcategories and Tasks
  3. 3.
    Clinical Note Generation 3
    • Definition
    • Subcategories and Tasks
  4. 4.
    Patient Communication and Education 5
    • Definition
    • Subcategories and Tasks
  5. 5.
    Medical Research Assistance 7
    • Definition
    • Subcategories and Tasks
  6. 6.
    Administration and Workflow 9
    • Definition
    • Subcategories and Tasks
Overview
  • Total Categories: 5
  • Entities per Category:
    • Definition: A description of the category.
    • Inclusion criteria: Guidelines determining which subcategories (and tasks) belong in the category.
    • Task performer: Individuals responsible for executing tasks under the category.
    • Subcategories: Each category contains multiple subcategories.
    • Tasks: Each subcategory comprises several tasks.
        Category
        - Definition
        - Inclusion criteria
        - Task performer
        - Sub category
            - Task
            - Task
            - Task
        - Sub category
            - Task
            - Task
            - Task
1. Clinical Decision Support
Definition: Analyzing patient-specific data to provide evidence-based recommendations to clinicians.
Inclusion criteria: Actions generating actionable insights based on patient data and medical evidence.
Task performer: Healthcare professionals (clinicians, nurses, and other practitioners).
Subcategories and Tasks:
  • Supporting Diagnostic Decisions
    • Recognize disease patterns from symptoms/vitals
    • Interpret diagnostic tests (ECG, spirometry)
    • Generate follow-up questions
    • Generate differential diagnoses
    • Interpret lab results
    • Detect image findings
    • Perform medical calculations
    • Evaluate social determinants of health
    • Track lab trends
    • Process intake information
  • Planning Treatments
    • Check drug interactions
    • Match protocols / screen contraindications
    • Suggest clinical pathways
    • Predict treatment response
    • Make collaborative decisions
    • Evaluate treatment accessibility
  • Predicting Risks and Outcomes
    • Predict deterioration, readmission, disease progression
    • Predict outcomes, adverse events, discharge readiness
    • Predict need for procedures or referrals
    • Manage preventive screening
  • Providing Clinical Knowledge Support
    • Apply guidelines and best practices
    • Answer medical knowledge questions
    • Track protocol compliance
    • Assess care quality
2. Clinical Note Generation
Definition: Creating structured records of patient care.
Inclusion criteria: Actions producing or modifying official clinical records.
Task performer: Providers, scribes, and documentation specialists.
Subcategories and Tasks:
  • Documenting Patient Visits
    • Generate progress, consultation, ED, admission, and discharge notes
    • Synthesize external/internal records
    • Summarize clinical documents
    • Generate team assessments
  • Recording Procedures
    • Generate OR, bedside, specialized procedure notes
  • Documenting Diagnostic Reports
    • Generate imaging, pathology, test, and genomic reports
  • Documenting Care Plans
    • Document treatment plans, care protocols, nursing plans, advance planning
3. Patient Communication and Education
Definition: Transmitting health information to enable patient understanding.
Inclusion criteria: Acts that support informed patient participation.
Task performer: Providers, coordinators, educators.
Subcategories and Tasks:
  • Providing Education Resources
    • Simplify disease info, risk factors, treatment
    • Explain insurance and billing
  • Delivering Personalized Instructions
    • Generate medication, procedure, and home care instructions
    • Explain follow-up and recovery
  • Patient-Provider Messaging
    • Triage messages, analyze symptoms
    • Handle refills, appointments, questions
    • Share results, draft responses
  • Enhancing Accessibility
    • Generate visual aids, translate content, make content accessible
  • Facilitating Engagement
    • Send reminders, preventive care, track goals
    • Collect feedback, support counseling and groups
4. Medical Research Assistance
Definition: Analyzing clinical data and literature to advance medical knowledge.
Inclusion criteria: Actions transforming data into scientific evidence.
Task performer: Researchers and epidemiologists.
Subcategories and Tasks:
  • Conducting Literature Research
    • Screen reviews, summarize papers
    • Analyze citations, synthesize evidence, identify gaps
  • Analyzing Research Data
    • Analyze trials, compare treatments, assess outcomes
    • Conduct cohort studies
  • Recording Research Processes
    • Support protocols, grants, manuscripts, statistics
  • Ensuring Research Quality
    • Validate methods, assess bias, handle regulatory needs
  • Managing Enrollment
    • Screen and match patients, track and document recruitment
5. Administration and Workflow
Definition: Orchestrating clinical operations from scheduling to billing.
Inclusion criteria: Actions managing logistics, resources, and flow.
Task performer: Administrators, staff, billing specialists.
Subcategories and Tasks:
  • Scheduling Resources and Staff
    • Schedule staff, manage inventory, equipment, and facilities
    • Monitor institutional metrics
  • Overseeing Financial Activities
    • Generate/document billing, insurer communication
    • Analyze revenue/costs, estimate out-of-pocket costs
  • Organizing Workflow Processes
    • Schedule appointments, process referrals/documents
    • Handle information requests
  • Care Coordination and Planning
    • Evaluate admissions, coordinate providers
    • Manage post-discharge planning and transitions
Appendix B Clinician Participant Demographics
We validated our task taxonomy through a survey of 29 clinicians representing 14 medical specialties across 4 institutions. Tables 6 and 5 provide detailed breakdowns of participant specialties and affiliations.
Affiliation Number of Participants
Stanford University 26
Mayo Clinic 1
Oregon Health & Science University 1
Flourish Research 1
Table 5: Number of participants per affiliation.
Speciality Number of Participants
Internal Medicine 9
Radiology 2
Otolaryngology 2
Neurology 2
Family Medicine 2
Dermatology 2
Anesthesiology 2
Radiation Oncology 2
Emergency Medicine 1
Ophthalmology 1
Psychiatry 1
Pathology 1
Pediatrics 1
Nuclear Medicine 1
Table 6: Number of participants per speciality.
Appendix C List of Benchmarks in MedHELM
Category Dataset Name Dataset Description Access Level Curation Status
Clinical Decision Support MedCalc-Bench A dataset which consists of a patient note, a question requesting to compute a specific medical value, and a ground truth answer. Public Existing
CLEAR A dataset that evaluates medical condition detection from patient notes using yes/no/maybe classifications. Private Formulated
MTSamples A dataset that provides transcribed medical reports and prompts models to generate appropriate treatment plans. Public Formulated
Medec A dataset containing medical narratives with error detection and correction pairs. Public Existing
EHRSHOT A dataset given a patient record of EHR codes, classifying if an event will occur at a future date or not. Gated Existing
HeadQA A collection of biomedical multiple-choice questions for testing medical knowledge. Public Existing
Medbullets A USMLE-style medical question dataset with multiple-choice answers and explanations. Public Existing
MedAlign A dataset that asks models to answer questions/follow instructions over longitudinal EHR. Gated Existing
ADHD-Behavior A dataset that classifies whether a clinical note contains a clinician recommendation for parent training in behavior management, which is the first-line evidence-based treatment for young children with ADHD. Private New
ADHD-MedEffects A dataset that classifies whether a clinical note contains documentation of side effect monitoring (recording of absence or presence of medication side effects), as recommended in clinical practice guidelines. Private New
Clinical Note Generation DischargeMe DischargeMe is a discharge instruction generation dataset and brief hospital course generation dataset collected from MIMIC-IV data, considering only the discharge text as well as the radiology report text. Gated Existing
ACI-Bench A dataset of patient-doctor conversations paired with structured clinical notes. Public Existing
MTSamples Procedures A dataset that provides a patient note regarding an operation, with the objective to document the procedure. Public Formulated
MIMIC-RRS A dataset containing radiology reports with findings sections from MIMIC-III paired with their corresponding impression sections, used for generating radiology report summaries. Gated Formulated
MIMIC-BHC A summarization task using a curated collection of preprocessed discharge notes paired with their corresponding brief hospital course (BHC) summaries. Gated Existing
NoteExtract A dataset containing free form text of a clinical health worker care plan, with the associated goal being to restructure that text into a given format. Private New
Patient Communication and Education MedicationQA Open text question-answer pairs regarding consumer health questions about medication. Public Existing
PatientInstruct A dataset containing case details used to generate customized post-procedure patient instructions. Private New
MedDialog A collection of doctor-patient conversations with corresponding summaries. Public Existing
MedConfInfo A dataset of clinical notes from adolescent patients used to identify sensitive protected health information that should be restricted from parental access. Private New
MEDIQA-QA A dataset including a medical question, a set of candidate answers,relevance annotations for ranking, and additional context to evaluate understandingand retrieval capabilities in a medical setting. Public Existing
MentalHealth A dataset containing a counselor and mental health patient conversation, where the objective is to generate an empathetic counselor response. Private New
PrivacyDetection A dataset that determines if a message leaks any confidential information from the patient Private New
ProxySender A dataset that determines if a message was sent by a proxy user Private New
Medical Research Assistance PubMedQA A dataset that provides pubmed abstracts and asks associated questions yes/no/maybe questions. Public Existing
EHRSQL Given a natural language instruction, generate an SQL query that would be used in clinical research. Public Existing
BMT-Status A dataset containing patient notes with associated questions and answers related to bone marrow transplantation. Private New
RaceBias A collection of LLM outputs in response to medical questions with race-based biases, with the objective being to classify whether the output contains racially biased content. Public Formulated
N2C2-CT A dataset that provides clinical notes and asks the model to classify whether the patient is a valid candidate for a provided clinical trial. Gated Existing
MedHallu A dataset of PubMed articles and associated questions, with the objective being to classify whether the answer is factual or hallucinated. Public Existing
Administration and Workflow HospiceReferral A dataset evaluating performance in identifying appropriate patient referrals to hospice care. Private New
MIMIC-IV Billing Code A dataset pairing clinical notes from MIMIC-IV with corresponding ICD-10 billing codes. Gated Existing
ClinicReferral A dataset containing manually curated answers to questions regarding patient referrals to the Sequoia clinic. Private New
CDI-QA A dataset built from Clinical Document Integrity (CDI) notes, to assess the ability to answer verification questions from previous notes. Private New
ENT-Referral A dataset designed to evaluate performance in identifying appropriate patient referrals to Ear, Nose, and Throat specialists. Private New
Table 7: Overview of all 35 benchmarks included in MedHELM, categorized by category, access level and curation status.
Appendix D LLM-based filtering
We assessed the quality of gold standard responses for reformulated benchmarks. Using GPT-4o-mini as a classifier, we identified reasonable versus unreasonable gold standard responses with the following prompt: Evaluate the following text: {prompt}{context} against the following reference: {gold standard response}. Return a score of 0 if the gold standard response appears to be using outside information that is not present in the input text and is not expected to be external knowledge known by a clinician, and thus is an unreasonable gold standard response to the question. Give a score of 1 if the reference is a reasonable response to the prompt. Only give a score of 0 if the reference does not make sense with the prompt or the question couldn’t be answered in this way without additional information. Only return the score, no other text.
As shown in Figure  6, 3 out of 5 reformulated benchmarks have less than 5% detected error in the gold standard response. For the remaining 2 benchmarks (MIMIC-RRS and MTSamples-Procedures), our evaluation methodology proved robust to gold standard response imperfections, with consistent mean scores between flagged and accepted answers. This resilience stems from our LLM-jury prompt design, which instructs judges to reference gold standards only when necessary.
Figure 6: We filter the question-answer pairs into a positive and negative group via LLM-based filtering. We see that the mean score of both groups remains relatively consistent, indicating stability in judging.
Appendix E Minimum Detectable Effect Evaluations
We calculate the minimum detectable effect at the benchmark level with the following formulation, where b is the benchmark, ij are two models to compare in a paired evaluation selected from all 9 models in $M$, $\sigma_{b}^{ij}$ is the standard deviation of the difference between the two selected model outputs for every question in the benchmark, and $n_{b}$ is the number of questions in the benchmark. $\mathrm{SD}_{b}$ is the standard deviation over the total pairwise MDE scores for the given benchmark, $b$. Additionally, we set $\alpha=0.05$ and $\beta=0.20$
$\mathrm{MDE}_{b}=\frac{1}{|\mathcal{M}|}\sum_{(i,j)\in\mathcal{M}}\frac{(z_{1-\frac{\alpha}{2}}+z_{1-\beta})\cdot\sigma_{b}^{(i,j)}}{\sqrt{n_{b}}}\quad\pm\quad\mathrm{SD}_{b}$
Table 8 shows the MDE for each benchmark and supports that the differences identified between model performance are significant.
Benchmark MDE
PatientInstruct $0.008\pm 0.002$
MIMIC-RRS $0.018\pm 0.005$
CLEAR $0.054\pm 0.003$
ENT-Referral $0.045\pm 0.005$
NoteExtract $0.009\pm 0.011$
CDI-QA $0.050\pm 0.007$
N2C2 $0.087\pm 0.010$
MentalHealth $0.032\pm 0.011$
PrivacyDetection $0.040\pm 0.013$
ProxySender $0.074\pm 0.009$
MIMIC-IV-BHC $0.019\pm 0.006$
ADHD-Behavior $0.056\pm 0.008$
ClinicReferral $0.050\pm 0.011$
DischargeMe $0.008\pm 0.002$
MedConfInfo $0.031\pm 0.004$
EHRSHOT $0.033\pm 0.004$
MIMIC-IV Billing Code $0.015\pm 0.003$
BMT-Status $0.079\pm 0.009$
HospiceReferral $0.042\pm 0.005$
ADHD-MedEffects $0.032\pm 0.009$
MedAlign $0.041\pm 0.005$
EHRSQL $0.034\pm 0.009$
MTSamples Procedures $0.020\pm 0.010$
ACI-Bench $0.012\pm 0.002$
MedicationQA $0.012\pm 0.004$
MedCalc-Bench $0.037\pm 0.005$
MEDIQA $0.018\pm 0.008$
MedBullets $0.079\pm 0.006$
MedDialog $0.007\pm 0.002$
MTSamples $0.011\pm 0.002$
Medec $0.052\pm 0.011$
PubMedQA $0.037\pm 0.004$
HeadQA $0.030\pm 0.007$
MedHallu $0.030\pm 0.004$
RaceBias $0.092\pm 0.016$
Table 8: Minimum Detectable Effect (MDE) values for various benchmarks
Appendix F LLM-Jury Prompt
While each jury prompt is customized for the given benchmark, we include the ACI-Bench jury prompt to serve as an example of how our LLM-Jury prompt is structured.
You are a medical expert tasked with evaluating the quality of a
generated clinical note based on a doctor-patient conversation.
Your goal is to assess how well the note captures the clinical information from the conversation and
compare it to the reference note (gold standard) in terms of accruacy, completeness and clarity.
The conversation will be provided in these tags:
<conversation>{{QUESTION}}</conversation>
The generated note will be provided in these tags:
<response>{{RESPONSE}}</response>
The reference note will be provided in these tags:
<gold_response>{{GOLD_RESPONSE}}</gold_response>
Carefully review the <response> based on the <conversation> and compare it to the <gold_response> when needed.
For each of the following criteria, rate the response on a scale of 1 to 5 (1 = very poor, 5 = excellent),
and provide a short justification for your score.
Evaluation Criteria:
Accuracy (1-5) - Does the note provide correct clinical information based on the conversation?
Completeness (1-5) - Does the note include all important medical details from the conversation?
Clarity (1-5) - Is the note written clearly and organized in a standard clinical format for clinicians
Output Format:
Output the evaluation as a single valid JSON object matching the following structure:
{"accuracy":{"score":0,"explanation":"Explain why this score was given."},
"completeness":{"score":0,"explanation":"Explain why this score was given."},
"clarity":{"score":0,"explanation":"Explain why this score was given."}}
Ensure the output is valid JSON:
- Use **double quotes** (") for all keys and string values.
- When quoting text or sections inside the explanations, use escaped double quotes (\") to
maintain valid JSON formatting.
- Do not include any additional information in the output.
Appendix G LLM Inference Costs
Table 9 summarizes the per-token inference costs of the LLMs used in MedHELM as of 05/12/2025.
Model Release Date Cost (per million tokens)
Input Tokens Output Tokens
GPT-4o 05/13/2024 \$2.50 \$10.00
GPT-4o Mini 07/18/2024 \$0.15 \$0.60
Llama 3.3 Instruct (70B) 12/06/2024 \$0.71 \$0.71
Gemini 2.0 Flash 02/01/2025 \$0.15 \$0.60
Gemini 1.5 Pro (001) 05/24/2024 \$1.25 \$5.00
Claude 3.5 Sonnet 10/22/2024 \$3.00 \$15.00
Claude 3.7 Sonnet 02/19/2025 \$3.00 \$15.00
o3-mini 01/31/2025 \$1.20 \$4.84
DeepSeek R1 01/20/2025 \$1.20 \$4.84
Table 9: Per-token inference costs of LLMs used in MedHELM.